Given f(x)=2x+1/3x-2. Find the value(s) of k for which f(2k-4) is undefined. Thanks for your help.
Answer by TimothyLamb(4379)

assume the equation has the following form:
if this form is incorrect, the answer will be incorrect:
---
f(x) = 2x + 1/(3x) - 2
---
f(2k-4) = 2(2k-4) + 1/(3(2k-4)) - 2
f(2k-4) = 4k - 8 + 1/(6k-12) - 2
f(2k-4) = 4k + 1/(6k-12) - 10
---
f(2k-4) is undefined when 1/(6k-12) is undefined:
---
1/(6k-12) is undefined when (6k-12)=0
---
6k - 12 = 0
6k = 12
k = 12/6
k = 2
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---